OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} {\frac{ 2^{2x} }{ 5^{x-1}}}\]

OpenStudy (anonymous):

\[2^{2x}=(2^2)^x=4^x\] might be a good start

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

I'm thinking. I don't have a good way to deal with the exponents though.

OpenStudy (anonymous):

make them equal

OpenStudy (anonymous):

multiply by \(\frac{1}{5}\) and they will be equal or if you think that is illegal, multipy top and bottom by 5, pull the top 5 out in front of the limit

OpenStudy (anonymous):

so then I have 5* lim 4^n/5^n which would be 0

OpenStudy (anonymous):

since 5^x is growing faster

OpenStudy (anonymous):

yeah if you want to be real slick you could say \[5\lim_{x\to \infty}\left(\frac{4}{5}\right)^x=0\]

OpenStudy (anonymous):

ok,cool. Thanks!

OpenStudy (anonymous):

yw